OpenStudy (anonymous):

dP/dt = p(2-P) , p(0)=1/2 .... How do I solve for P(t) for the initial value?

terenzreignz (terenzreignz):

It looks like it's more-or-less separable. We get. \[\Large \frac{dP}{P(2-P)}=dt\]

terenzreignz (terenzreignz):

Can you integrate both sides?

OpenStudy (anonymous):

int(1/p(2-p))dp = 1 dt then int by parts a/p +b/2-p =int1/2p + int1/2(2-p) =1/2(lnp) - 1/2

OpenStudy (anonymous):

1/2(lnp) - 1/2(ln(2-p) = t +C

OpenStudy (anonymous):

then, do I solve for C? from here I'm not sure what to do...

terenzreignz (terenzreignz):

Something's amiss... are you sure it isn't \[\Large \frac1{2}\ln(P) - \frac12\ln(P-2)\] ?

OpenStudy (anonymous):

well, after integration by parts i get \[1/2 \int\limits(1/p) + 1/2 \int\limits(1/2-p))\] then after integrate shouldnt that give \[1/2(\ln(p)) - 1/2(\ln(2-p))\]?

OpenStudy (anonymous):

the difference I see is in the ln(p-2) vs ln(2-p)

terenzreignz (terenzreignz):

Hang on... \[\Large \frac12\int\frac1p \color{green}{dp} + \frac12\int \frac1{2-p}\color{green}{dp}\] right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

First one's no problem: \[\Large \frac12\ln(p) + \frac12\int\frac1{2-p}\color{green}{dp}\]

terenzreignz (terenzreignz):

Hang on...

terenzreignz (terenzreignz):

Wait, I see the problem... we're both right... I just forgot that the actual integral of 1/x is \[\Large \int \frac1xdx = \ln|x|+C\] So it doesn't matter really if it were \[\Large \ln(P-2) \] or \[\Large \ln(2-P)\] LOL moving on...

terenzreignz (terenzreignz):

But just to be sure, let's have it as \[\Large \frac12\ln|P|-\frac12\ln|P-2| = t+C \]

OpenStudy (anonymous):

lol, okay so assuming that correct I think I solve doe C using the initial values p(0)=1/2 then splve for P and plug in the C value... Is that the right idea?

terenzreignz (terenzreignz):

Yup. Replace all t with 0 and all p with 1/2

OpenStudy (anonymous):

okay.. I get C = (ln(3/4))/2

terenzreignz (terenzreignz):

I think the arithmetic was a bit off...

OpenStudy (anonymous):

Okay I need to crunch away on this for a while, as long as I know the process is correct I can do it. Thanks so much for your help!